# LLMs Can’t Plan, But Can Help Planning in LLM-Modulo Frameworks

### 06 Feb 2024 v2

@@ Abstract
There is considerable confusion about the role of Large Language Models (LLMs) in planning and reasoning tasks. On one side are over-optimistic claims that LLMs can indeed do these tasks with just the right prompting or self-verification strategies. On the other side are perhaps over-pessimistic claims that all that LLMs are good for in planning/reasoning tasks are as mere translators of the problem specification from one syntactic format to another, and ship the problem off to external symbolic solvers. In this position paper, we take the view that both these extremes are misguided. We argue that auto-regressive LLMs cannot, by themselves, do planning or self-verification (which is after all a form of reasoning), and shed some light on the reasons for misunderstandings in the literature. We will also argue that LLMs should be viewed as universal approximate knowledge sources that have much more meaningful roles to play in planning/reasoning tasks beyond simple front-end/back-end format translators. We present a vision of LLM-Modulo Frameworks that combine the strengths of LLMs with external model-based verifiers in a tighter bi-directional interaction regime. We will show how the models driving the external verifiers themselves can be acquired with the help of LLMs. We will also argue that rather than simply pipelining LLMs and symbolic components, this LLM-Modulo Framework provides a better neuro-symbolic approach that offers tighter integration between LLMs and symbolic components, and allows extending the scope of model-based planning/reasoning regimes towards more flexible knowledge, problem and preference specifications.
@@ 1 Introduction
Large Language Models (LLMs), essentially n-gram models on steroids which have been trained on web-scale language corpora (or, effectively, our collective consciousness), have caught the imagination of the AI research community with linguistic behaviors that no one expected text completion systems to possess. Their seeming versatility has led many researchers to wonder whether they can also do well on planning and reasoning tasks typically associated with System 2 competency. On the face of it, this doesn’t seem to ring true, as both by training and operation, LLMs are best seen as a giant pseudo System 1 Kahneman (2011) (see Figure 1). Even from a pure engineering perspective, a system that takes constant time to produce the next token cannot possibly be doing principled reasoning on its own. [^1] Think of asking an LLM an yes/no question–is this theorem logically entailed by this first-order logic knowledge-base. This is well-known to be a semi-decidable problem. Ask yourself if the LLM will take longer in answering the question. (If you are thinking Chain-of-thought prompts or training with step-by-step data, consider that you are essentially changing the nature of the original prompt/training). Not surprisingly, initial excitement based on anecdotal performance of LLMs on reasoning tasks Bubeck et al. (2023) has dissipated to some extent by the recent spate of studies questioning the robustness of such behaviors–be it planning Valmeekam et al. (2023c); Kambhampati (2023), simple arithmetic and logic Dziri et al. (2023), theory of mind abilities Ullman (2023); Verma et al. (2024), or general mathematical and abstract benchmarks McCoy et al. (2023); Gendron et al. (2023). Despite this, a steady stream of claims continue to be made in the literature about the planning and reasoning capabilities of LLMs.
@@
In an ironic juxtaposition to this unwarranted optimism about the planning and reasoning abilities of LLMs, there is also unwarranted pessimism about the roles LLMs can play in planning/reasoning tasks. Several efforts (e.g. Liu et al. (2023); Pan et al. (2023); Xie et al. (2023)) advocate using LLMs only as glorified translators–for converting reasoning problems embedded in textual format to symbolic representations, which are then pawed off to external classical symbolic solvers (with all their attendant expressivity and search complexity challenges Doyle & Patil (1991)). [^2] In some circles, this unidirectional pipeline has come to be given the undeserved badge of neuro-symbolic architecture.
@@
Figure 1: An informal account of viewing LLM as a giant external non-veridical memory that acts as a pseudo System 1
@@
In truth, LLMs are a whole lot more than machine translators. They are a kind of approximate knowledge sources (albeit sans guarantees) trained on our collective consciousness. While it is unlikely that they will have System 2 competencies by themselves, they can nevertheless be valuable resources in solving System 2 tasks. To put it another way, the problem with Alchemy of yore is not that Chemistry is useless, but that people wanted to delude themselves that Chemistry–a pretty amazing discipline on its own merits–can be Nuclear Physics if you prompt it just so. The confusions regarding LLM abilities, or should we say, LLM alchemy, seems to be not that much different–oscillating between ignoring what they are good at, and ascribing them abilities they don’t have.
@@
The goal of this position paper is to introduce some clarity into this confusing state of affairs oscillating between over-optimism and over-pessimism. Simply put, we take the stance that LLMs are amazing giant external non-veridical memories that can serve as powerful cognitive orthotics for human or machine agents, if rightly used. The underlying n-gram nature makes them effortlessly intermix what would be considered disparate fields of study (not surprisingly, LLMs are seen to be very good at making/finding analogies!) The challenge is to leverage them without wrongly ascribing to them capabilities they don’t possess. The LLM-Modulo framework proposed in this position paper tackles this challenge.
@@
For the sake of concreteness, we focus on planning tasks, especially as studied in the automated planning community Ghallab et al. (2004). The central position of the paper is that LLMs cannot plan themselves but can play a variety of constructive roles in solving planning tasks–especially as approximate knowledge sources and candidate plan generators in the so-called LLM-Modulo Frameworks in conjunction with external sound model-based verifiers.
@@
We support this position by first reviewing literature that establishes that LLMs cannot be used as planners or plan verifiers themselves (Section 2). We also discuss why there are claims about planning/verification abilities in the first place, in the process hopefully clarifying some prevalent misunderstandings.
@@
Second, we will propose a framework that allows us to leverage LLMs effectively in planning tasks, by combining them with external critics, verifiers and humans. We call this an LLM-Modulo Framework (a name loosely inspired by SAT Modulo Theories Nieuwenhuis & Oliveras (2006)); see Figure 3. LLMs play a spectrum of roles in this architecture, from guessing candidate plans, to translating those plans into syntactic forms that are more accessible to external critics, to helping end users flesh out incomplete specifications, to helping expert users acquire domain models (that in turn drive model-based critics). All this leveraging of LLMs is done without ascribing to them any planning or verification abilities. The LLM ideas are vetted by external critics, thus ensuring that the plans generated in this architecture can have formal correctness guarantees where possible.
One question regarding this LLM-Modulo architecture from a planning perspective is whether it is more than a gratuitous attempt to shoe-horn LLMs to solve planning problems, when there are more formal combinatorial planning systems available in multiple communities Ghallab et al. (2004). Compared to a planner that is guaranteed to be correct in a narrow set of domains, LLMs may likely be good at generating plausible (but not guaranteed to be correct) plan heuristics/suggestions in many more scenarios. Thus, unlike the traditional planning architectures studied in AI Ghallab et al. (2004), which put a priori constraints on the expressiveness of the problems that can be posed to the planner (to wit, the different expressiveness levels of the PDDL specification McDermott et al. (1998)), the LLM-Modulo architecture puts no such restrictions. In this sense, it is more representative of real-world planning problems such as those in NASA mission planning, where the different critics–human and automated–are at best able to give “no objection” certificates, clearing it from their perspective. (Indeed, both deep space network planning and mars rover task planning are done via a collective human blackboard. Johnston et al. (2014) Bresina et al. (2004).) Note that this is starkly different from just sending an unvetted plan out to execution (as would be the case if we have LLMs operate in autonomous mode to guess plans). Generalizing planning and reasoning frameworks this way is consistent with the Doyle & Patil’s call to the Knowledge Representation community of yore Doyle & Patil (1991).
@@ 2 Planning-centered Limitations of LLMs
In this section, we will first review the literature that calls into question the claims about planning and self-verification capabilities of LLMs. Subsequently, we will also provide some possible reasons for the claims to the contrary made in the literature.
@@ 2.1 LLMs cannot generate executable plans in autonomous mode
Despite the initial claims about the planning capabilities of LLMs Bairi et al. (2023); Yao et al. (2023b); Shinn et al. (2023); Huang et al. (2022); Hao et al. (2023) several recent studies independently confirm that LLMs are not actually able to generate executable plans when they are used in autonomous modes Valmeekam et al. (2023c); Liu et al. (2023); Silver et al. (2022). For example, in Valmeekam et al. (2023c, b), the authors evaluate LLMs ability to generate correct plans on a suite of planning problem instances based on the kinds of domains employed in the International Planning Competition IPC (1998). To eliminate the subjective aspect of analysis that forms the core part of many earlier efforts on evaluating the reasoning capabilities of LLMs, they automate the evaluation by leveraging models and tools from the automated planning community.
They show that the results in the autonomous mode are pretty bleak. On average, only about 12% of the plans that the best LLM (GPT-4) generates are actually executable without errors and reach their goals. They show that the choice of the specific LLM (they have tested the family of GPT LLMs including GPT-4 OpenAI (2023), GPT-3.5 OpenAI (2022), InstructGPT-3 Ouyang et al. (2022) and GPT-3 Brown et al. (2020)). They also sho that fine-tuning does not seem to have a major effect on this dismal performance. They also show that the performance deteriorates further if the names of the actions and objects in the domain are obfuscated–a change that doesn’t in any way affect the performance of the standard AI planners. This latter further suggests that LLMs are more likely doing approximate retrieval of plans than planning.
@@ 2.2 LLMs cannot verify plans and thus cannot improve by self-critiquing
There still exists considerable optimism that even if LLMs can’t generate correct solutions in one go, their accuracy might improve in an iterative prompting regime, where LLMs will be able to “self-critique” their candidate solutions and refine them to the point of correctness Yao et al. (2023b, a); Shinn et al. (2023); Weng et al. (2023); Huang et al. (2022). This belief seems to rest largely on the assumption that verification of correctness should be easier than generation for many reasoning problems–a rather classical argument from computational complexity. There are grounds to be skeptical of this assumption as the complexity of the reasoning task should be irrelevant to LLM performance if what they are doing is approximate retrieval. In general, unless LLMs are trained not just on “correct data,” but also “corrections data,” there is no a priori reason to believe that their critiques would even be approximately relevant, let alone actually correct.
@@
Two recent studies–one on plan verification Valmeekam et al. (2023a) and the other on CSP verification Stechly et al. (2023) seem to throw cold water on this optimism. In Stechly et al. (2023), the authors systematically investigate the effectiveness of iterative prompting in the context of Graph Coloring, a canonical NP-complete reasoning problem. Their methodology involves a principled empirical study of the performance of GPT4 on two tasks: solving a large suite of random graph coloring instances and, separately, verifying the correctness of the candidate colorings–both in direct (i.e., return the first solution generated by the LLM) and iterative modes. In iterative modes, they experiment both with an LLM critiquing LLM-produced solutions and an external, guaranteed correct reasoner verifying solutions. In both cases, they analyze whether the content of criticisms actually affects bottom-line performance.
@@
Their results indicate that in direct mode, LLMs are, perhaps not surprisingly, pretty bad at solving graph coloring instances. More interestingly, they are no better at verifying solutions. In iterative modes, given the inability of LLMs to verify solutions, it should come as no surprise that their experiments also show that the strategy of LLMs self-critiquing their solutions does not improve over the baseline. They report that the perforance is in fact worse because the system can’t recognize a correct coloring and thus merrily passes over fortuitously correct colorings it has generated, ending up with a wrong one! Similar results have also been reported for planning problems in Valmeekam et al. (2023c).
@@
One important corollary of the fact that LLMs cannot self-critique their plans is that they can’t also self-improve by generating synthetic data by generating plans themselves, critiquing the plans by themselves to improve them, and then using those to fine-tune themselves, as has been claimed in the literature Wang et al. (2022) [^3] Contrary to the claim of “self-improvement”, works like Wang et al. (2022) actually heavily depend on external knowledge (crafted seed examples) and critics (filtering step). Huang et al. (2023b); see Section 3.3.
@@ 2.3 Analyzing Claims to the Contrary in the Literature
Given that LLMs can neither guarantee correct generation nor correct verification of plans, as discussed in the previous sections, one obvious question is why the literature is replete with claims contrary to this Bairi et al. (2023); Yao et al. (2023b); Shinn et al. (2023); Yao et al. (2023a); Weng et al. (2023); Huang et al. (2022).
@@
Claims about Planning: To analyze the planning claims, we need to first understand that solving planning tasks requires (a) having the necessary planning domain knowledge–the actions and their preconditions, effects; the standard hierarchical recipes (e.g. task reduction schemas in HTN planning), past cases/plans, etc., and (b) being able to assemble this planning knowledge into an executable plan that takes care of any subgoal/resource interactions. The first can be called the knowledge acquisition and the second reasoning/planning part. Many of the papers claiming planning abilities of LLMs, on closer examination, wind up confusing general planning knowledge extracted from the LLMs for executable plans. When all we are looking for are abstract plans, such as “wedding plans,” with no intention of actually executing the said plans, it is easy to confuse them for complete executable plans. Indeed, our close examination of several works claiming planning capabilities for LLMs Kambhampati et al. (2023) suggests that they either work in domains/tasks where subgoal interactions can be safely ignored Yao et al. (2023b); Shinn et al. (2023) [^4] Although domains like AlfWorld Shridhar et al. (2021) do have sub-goal interactions for successful task completion, Yao et al. (2023b) and Shinn et al. (2023) ignore these interactions relying on the ergodic nature of the domain when prompting LLMs for generating plans. –either because they are just working on a single goal, or because the world is forgiving and ergodic; or delegate the interaction resolution (reasoning) to the humans in the loop (who, through repeated prompting, have to “correct” the plan). Sometimes, in common sense domains, or with enough fine-tuning, the “assembling” part may also be obviated by having seen a case that pretty much corresponds to the problem that needs to be solved. Not surprisingly, the work by Valmeekam et al. (2023c) shows that if the action interactions are removed by relaxing the world models, then the ability of LLMs to guess executable plans improves. Without these assumptions or mitigations, the plans that come out of LLMs may look reasonable to the lay user, and yet lead to execution time interactions and errors. [^5] These issues are illustrated in part by a recent news story Kugel & Hiltner (2023) about the proliferation of travel planning books, mostly auto-extracted from LLMs, and the ensuing disappointment of the unsuspecting end users who buy them mistaking them for usable plans!
The fact that LLMs are often good at extracting planning knowledge can indeed be gainfully leveraged. As shown in recent works Guan et al. (2023), LLMs can be a rich source of approximate models of world/domain dynamics and user preferences, as long as the humans (and any specialized critics) in the loop verify and refine those models, and give them over to model-based solvers. This way of using LLMs has the advantage that the humans need only be present when the dynamics/preference model is being teased out and refined, and the actual planning after that can be left to sounder planning frameworks with correctness guarantees, such as LLM-Modulo framework we propose.
@@
Such an overall approach has striking similarities to knowledge-based AI systems of yore, with LLMs effectively replacing the “knowledge engineer” (see Figure 2). Given the rather quixotic and dogmatic shift of AI away from approaches that accept domain knowledge from human experts that some writers termed “Polanyi’s Revenge” (c.f. Kambhampati (2021)), this new trend of using LLMs as knowledge sources can be viewed as a form of avenging Polanyi’s revenge! Indeed, LLMs make it easy to get problem-specific knowledge as long as we are willing to relax the correctness requirements of that knowledge. In contrast to the old knowledge engineering approaches, LLMs offer this without making it look like we are inconveniencing any specific human (we are, instead, just leveraging everything humans told each other on the Web!). So the million dollar question for reasoning tasks is: “how would you do robust planning if you have some doddering know-it-all ready to give you any kind of knowledge?” The LLM-Modulo Framework is a principled method for tackling this challenge.
@@
Figure 3: The proposed LLM-Modulo framework where LLMs act as idea generators and various external critiques that specialize in different aspects, critique the candidate plan.
@@
Claims about Self-Verification: Coming to the claims about LLM’s self-verification abilities, a closer look at the literature Yao et al. (2023a); Huang et al. (2023a) shows that those claims are either (i) made in the context of tacit knowledge tasks for which there is little possibility of a verifier (e.g. essay writing)–making it hard to evaluate whether LLM’s critiquing actually helped [^6] Paradoxically, the fact that it is infeasible to write sound verifiers for tacit knowledge tasks also makes it possible for everyone to be a critic. Think of R2 saying the paper could be made ”less dense” or the Peloton instructor critiquing Christopher Nolan film. or (ii) the external verification is carried out either by simulators Wang et al. (2023); Yao et al. (2023b) or simple calls to the underlying operating system (as is the case, for example, for the 24 puzzle in Yao et al. (2023a)).
@@
In a related vein, there is the recent Tree of Thoughts (ToT) paper Yao et al. (2023a), which has been pitched as a way to convert LLMs into some type of systematic search with self-verification. A closer look at the work however shows that ToT simply iteratively back-prompts the LLM until it comes up with a solution that is acceptable to an external verifier. Specifically, ToT employs a problem-specific prompt priming method. The “tree” in ToT is essentially a way to generate diverse priming prompts (that the authors set up in a problem specific way). In other words, despite the use of terminology of problem-solving agents Russell & Norvig (2010)–search tree, expansion etc., there is really no deeper connection to search-based agents. [^7] Our preliminary experiments also show that at least in 24 puzzle, a simple iterative prompting, even without a systematic prompt diversification, is quite competitive with the ToT framework.

### 12 Jun 2024 v3

# But Can Help Planning in LLM-Modulo Frameworks

@@ Abstract
We argue that auto-regressive LLMs cannot, by themselves, do planning or self-verification (which is after all a form of reasoning), and shed some light on the reasons for misunderstandings in the literature. We also argue that LLMs should be viewed as universal approximate knowledge sources that have much more meaningful roles to play in planning/reasoning tasks beyond simple front-end/back-end format translators. We present a vision of LLM-Modulo Frameworks that combines the strengths of LLMs with external model-based verifiers in a tighter bi-directional interaction regime. We will show how the models driving the external verifiers themselves can be acquired with the help of LLMs. We will also argue that rather than simply pipelining LLMs and symbolic components, this LLM-Modulo Framework provides a better neuro-symbolic approach that offers tighter integration between LLMs and symbolic components, extending the scope of model-based planning/reasoning regimes towards more flexible knowledge, problem and preference specifications.
LLMs, Planning, Reasoning
@@ 1 Introduction
Large Language Models (LLMs), essentially n-gram models on steroids which have been pre-trained on web-scale language corpora (or, effectively, our collective consciousness), have caught the imagination of the AI research community with linguistic capabilities that no one expected text completion systems to possess. Their seeming versatility has led many researchers to wonder whether they can also do well on planning and reasoning tasks typically associated with System 2 competency. On the face of it, this doesn’t seem to ring true, as both by training and operation, LLMs are best seen as a giant pseudo System 1 (Kahneman, 2011) (see Figure 1). Even from a pure engineering perspective, a system that takes constant time to produce the next token cannot possibly be doing principled reasoning on its own. [^1] Think of asking an LLM an yes/no question–is this theorem logically entailed by this first-order logic knowledge-base. This is well-known to be a semi-decidable problem. Ask yourself if the LLM will take longer in answering the question. (If you are thinking Chain-of-thought prompts or training with step-by-step data, consider that you are essentially changing the nature of the original prompt/training). Not surprisingly, initial excitement based on anecdotal performance of LLMs on reasoning tasks (Bubeck et al., 2023) has been dissipated to some extent by the recent spate of studies, including our own, questioning the robustness of such behaviors–be they planning (Valmeekam et al., 2023c; Kambhampati, 2024), simple arithmetic and logic (Dziri et al., 2023), theory of mind (Ullman, 2023; Verma et al., 2024b), or general mathematical and abstract benchmarks (McCoy et al., 2023; Gendron et al., 2023). Despite this, a steady stream of claims continue to be made in the literature about the planning and reasoning capabilities of LLMs. In light of questions about their planning capabilities, the head-long rush into agentic LLMs should be particularly concerning. After all, acting without the ability to plan is surely a recipe for unpleasant consequences!
@@
In an ironic juxtaposition to this unwarranted optimism about the planning and reasoning abilities of LLMs, there is also unwarranted pessimism about the roles LLMs can play in planning/reasoning tasks. Several efforts (e.g. (Liu et al., 2023; Pan et al., 2023; Xie et al., 2023)) advocate using LLMs only as glorified translators–converting reasoning problems embedded in textual format to symbolic representations, and pawning them off to external classical symbolic solvers (with all their attendant expressivity and search complexity challenges (Doyle & Patil, 1991)). [^2] In some circles, this unidirectional pipeline has been given the undeserved badge of neuro-symbolic architecture.
@@
Figure 1: An informal account of viewing an LLM as a giant external non-veridical memory that acts as a pseudo System 1
@@
In truth, LLMs can be a whole lot more than machine translators. They are a kind of approximate knowledge source (albeit sans guarantees) trained on our collective consciousness. While it is unlikely that they will have System 2 competencies by themselves, they can nevertheless be valuable resources in solving System 2 tasks. To put it another way, the problem with Alchemy of yore was not that Chemistry is useless, but that people wanted to delude themselves that Chemistry–a pretty amazing discipline on its own merits–can be Nuclear Physics if you prompt it just so. The confusions regarding LLM abilities, or should we say, LLM alchemy, doesn’t seem to be much different–oscillating between ignoring their strengths, and ascribing abilities they don’t have.
@@
The goal of this position paper is to introduce some clarity into this confusing state of affairs oscillating between over-optimism and over-pessimism. Simply put, we take the stance that LLMs are amazing giant external non-veridical memories that can serve as powerful cognitive orthotics for human or machine agents, if rightly used. The underlying n-gram nature makes them effortlessly intermix what would be considered disparate fields of study (not surprisingly, LLMs are seen to be very good at making/finding analogies!). The challenge is to leverage them without wrongly ascribing to them capabilities they don’t possess. The LLM-Modulo framework proposed in this position paper tackles this challenge.
@@
For the sake of concreteness, we consider planning tasks, especially as studied in the automated planning community (Ghallab et al., 2004). The central position of the paper is that LLMs cannot plan themselves but can play a variety of constructive roles in solving planning tasks–especially as approximate knowledge sources and candidate plan generators in so-called LLM-Modulo Frameworks, where they are used in conjunction with external sound model-based verifiers.
@@
We support this position by first reviewing literature, including our own works, that establishes that LLMs cannot be used as planners or plan verifiers themselves (Section 2). We also discuss why there are claims about planning/verification abilities in the first place, in the process hopefully clarifying some prevalent misunderstandings.
@@
Second, we will propose a framework that allows us to leverage LLMs effectively in planning tasks, by combining them with external critics, verifiers and humans. We call this an LLM-Modulo Framework (a name loosely inspired by SAT Modulo Theories (Nieuwenhuis & Oliveras, 2006)); see Figure 3. LLMs play a spectrum of roles in this architecture, from guessing candidate plans, to translating those plans into syntactic forms that are more accessible to external critics, to helping end users flesh out incomplete specifications, to helping expert users acquire domain models (that in turn drive model-based critics). All this leveraging of LLMs is done without ascribing to them any planning or verification abilities. The LLM ideas are vetted by external critics, thus ensuring that the plans generated in this architecture can have formal correctness guarantees where possible.
@@ 2 Planning-centered Limitations of LLMs
In this section, we will first review literature that calls into question claims about the planning and self-verification capabilities of LLMs. Subsequently, we will also provide some possible reasons for claims to the contrary made in the literature.
Domain Method Instances correct
GPT-4o GPT-4-Turbo Claude-3-Opus LLaMA-3 70B Gemini Pro GPT-4
Blocksworld (BW) One-shot 170/600 (28.33%) 138/600 (23%) 289/600 (48.17%) 76/600 (12.6%) 68/600 (11.3%) 206/600 (34.3%)
Zero-shot 213/600 (35.5%) 241/600 (40.1%) 356/600 (59.3%) 205/600 (34.16%) 3/600 (0.5%) 210/600 (34.6%)
Mystery BW (Deceptive) One-shot 5/600 (0.83%) 5/600 (0.83%) 8/600 (1.3%) 15/600 (2.5%) 2/500 (0.4%) 26/600 (4.3%)
Zero-shot 0/600 (0%) 1/600 (0.16%) 0/600 (0%) 0/600 (0%) (0/500) (0%) 1/600 (0.16%)
Table 1: Results of state-of-the-art LLMs GPT-4o, GPT-4-Turbo, Claude-3-Opus, Gemini Pro and LLaMA-3 70B for Plan Generation with prompts in natural language.
@@ 2.1 LLMs cannot generate executable plans in autonomous mode
Despite initial claims about the planning capabilities of LLMs (Bairi et al., 2023; Yao et al., 2023b; Shinn et al., 2023; Huang et al., 2022; Hao et al., 2023) several recent studies confirm that LLMs are not actually able to generate executable plans when they are used in autonomous modes (Valmeekam et al., 2023c; Liu et al., 2023; Silver et al., 2022). For example, in (Valmeekam et al., 2023c, b), we evaluate LLMs’ ability to generate correct plans on a suite of planning problem instances based on the kinds of domains employed in the International Planning Competition (IPC, 1998). To eliminate the subjective aspect of analysis that forms the core part of many earlier efforts to evaluate the reasoning capabilities of LLMs, we leverage models and tools from the automated planning community to automate evaluation.
We show that results in the autonomous mode are pretty bleak. On average, only about 12% of the plans that the best LLM (GPT-4) generates are actually executable without errors and goal-reaching. We show that the choice of LLM doesn’t have much bearing on this. We tested the family of GPT LLMs including GPT-4 (OpenAI, 2023), GPT-3.5 (OpenAI, 2022), InstructGPT-3 (Ouyang et al., 2022) and GPT-3 (Brown et al., 2020). We also show that fine-tuning does not seem to have a major effect on this dismal performance. We demonstrate that the performance deteriorates further if the names of the actions and objects in the domain are obfuscated–a change that doesn’t in any way affect the performance of the standard AI planners. This further suggests that LLMs are more likely doing approximate retrieval of plans than actual planning.
We continue to reconfirm these limitations over each of the more recently released LLMs, including Claude Opus, Gemini, GPT4-Turbo and GPT4-o. Table 1 shows that all the state of the art LLMs show dismal performance on PlanBench (Valmeekam et al., 2023b).
More recently, we have also investigated so-called “chain of thought” prompting (Stechly et al., 2024b), as well as ReAct-style step-by-step prompting (Verma et al., 2024a) and found that they too are largely ineffective in improving the planning performance of LLMs.
@@ 2.2 LLMs cannot verify plans and thus cannot improve by self-critiquing
There still exists considerable optimism that even if LLMs can’t generate correct solutions in one go, their accuracy might improve in an iterative prompting regime, where LLMs will be able to “self-critique” their candidate solutions and refine them to the point of correctness (Yao et al., 2023b, a; Shinn et al., 2023; Weng et al., 2023; Huang et al., 2022). This belief seems to rest largely on the assumption that verification of correctness should be easier than generation for many reasoning problems–a rather classical argument from computational complexity. There are grounds to be skeptical of this assumption as the complexity of the reasoning task should be irrelevant to LLM performance if what they are doing is approximate retrieval. In general, unless LLMs are trained not just on “correct data,” but also on “corrections data,” there is no a priori reason to believe that their critiques would even be approximately relevant, let alone actually correct.
@@
Two of our studies–one on plan verification (Valmeekam et al., 2023a) and the other on CSP verification (Stechly et al., 2023) seem to throw cold water on this optimism. In (Stechly et al., 2023), we systematically investigate the effectiveness of iterative prompting in the context of Graph Coloring, a canonical NP-complete reasoning problem. Our methodology involves a principled empirical study of the performance of GPT4 on two tasks: solving a large suite of random graph coloring instances and, separately, verifying the correctness of the candidate colorings–both in direct (i.e., return the first solution generated by the LLM) and iterative modes. In iterative modes, we experiment both with an LLM critiquing LLM-produced solutions and an external, guaranteed correct reasoner verifying solutions. In both cases, we analyze whether the content of criticisms actually affects bottom-line performance. A more recent paper further analyzes these results along with performance on the 24 puzzle–a task that has been used by some researchers claiming LLMs have the ability to self verify (Stechly et al., 2024a).
@@
Our results indicate that in direct mode, LLMs are, perhaps not surprisingly, pretty bad at solving graph coloring instances. More interestingly, they are no better at verifying solutions. In iterative modes, given the inability of LLMs to verify solutions, it should come as no surprise that our experiments also show that the strategy of LLMs self-critiquing their solutions does not improve over the baseline. We report that the performance is in fact worse because the system can’t recognize a correct coloring and thus merrily passes over fortuitously correct colorings it has generated, ending up with a wrong one! Similar results have also been reported for planning problems in (Valmeekam et al., 2023c).
@@
One important corollary of the fact that LLMs cannot self-critique their plans is that they also can’t self-improve by generating synthetic data, e.g. by generating plans themselves, critiquing the plans by themselves to improve them, and then using those to fine-tune themselves, as has been claimed in the literature (Huang et al., 2023b; Wang et al., 2022) [^3] Contrary to their claim of “self-improvement”, works like (Wang et al., 2022) actually heavily depend on external knowledge (crafted seed examples) and critics (filtering step)..
@@ 2.3 Analyzing Claims to the Contrary in the Literature
Given that LLMs can neither guarantee correct generation nor correct verification of plans, as discussed in the previous sections, one obvious question is why the literature is replete with claims contrary to this (Bairi et al., 2023; Yao et al., 2023b; Shinn et al., 2023; Yao et al., 2023a; Weng et al., 2023; Huang et al., 2022).
@@
Claims about Planning: To analyze planning claims, we need to first understand that solving planning tasks requires (a) having the necessary planning domain knowledge–the actions and their preconditions, effects; the standard hierarchical recipes (e.g. task reduction schemas in HTN planning), past cases/plans, etc., and (b) being able to assemble this planning knowledge into an executable plan that takes care of any subgoal/resource interactions. The first part can be called knowledge acquisition and the second reasoning/planning. On closer examination, many papers claiming LLMs have planning abilities wind up confusing general planning knowledge extracted from the LLMs for executable plans. When all we are looking for are abstract plans, such as “wedding plans,” with no intention of actually executing them, it is easy to confuse them for complete executable plans. Indeed, our close examination of several works claiming planning capabilities for LLMs (Kambhampati et al., 2023) suggests that they either work in domains/tasks where subgoal interactions can be safely ignored (Yao et al., 2023b; Shinn et al., 2023) [^4] Although domains like AlfWorld (Shridhar et al., 2021) do have sub-goal interactions for successful task completion, (Yao et al., 2023b) and (Shinn et al., 2023) largely ignore these interactions by either focusing on single subgoals or relying on the ergodic nature of the domain when prompting LLMs for generating plans (Verma et al., 2024a).–either because they are just working on a single subgoal, or because the world is forgiving and ergodic; or by delegating the interaction resolution (reasoning) to the humans in the loop (who, through repeated prompting, have to “correct” the plan). Sometimes, in common sense domains, or with enough fine-tuning, the “assembling” part may also be obviated by having seen a case that pretty much corresponds to the problem that needs to be solved. Not surprisingly, our work (Valmeekam et al., 2023c) shows that if the action interactions are removed by relaxing the world models, then the ability of LLMs to guess executable plans improves. Without these assumptions or mitigations, the plans that come out of LLMs may look reasonable to the lay user, and yet lead to execution time interactions and errors. [^5] These issues are illustrated in part by a recent news story (Kugel & Hiltner, 2023) about the proliferation of travel planning books, mostly auto-extracted from LLMs, and the ensuing disappointment of the unsuspecting end users who buy them mistaking them for usable plans!
@@
Figure 3: The proposed LLM-Modulo framework where LLMs act as idea generators and various external critics that specialize in different aspects, critique the candidate plan.
@@
Claims about Self-Verification: Coming to the claims about LLM’s self-verification abilities, a closer look at the literature (Yao et al., 2023a; Huang et al., 2023a) shows that those claims are either (i) made in the context of tacit knowledge tasks for which there is little possibility of a verifier (e.g. essay writing)–making it hard to evaluate whether LLM’s critiquing actually helped or (ii) the external verification is carried out either by simulators (Wang et al., 2023b; Yao et al., 2023b) or simple calls to the underlying operating system.
@@
In a related vein, there is the recent Tree of Thoughts (ToT) paper (Yao et al., 2023a), which has been pitched as a way to convert LLMs into some type of systematic search with self-verification. Specifically, ToT employs a problem-specific prompt priming method. The “tree” in ToT is essentially a way to generate diverse priming prompts (that the authors set up in a problem specific way). In other words, despite the use of terminology of problem-solving agents (Russell & Norvig, 2010)–search tree, expansion etc., there is really no deeper connection to search-based agents.